### v1

# LakeVilla: Multi-Table Transactions for Lakehouses

Abstract.
Data lakehouses (LHs) are at the core of current cloud analytics stacks by providing elastic, relational compute on data in cloud data lakes across vendors. For relational semantics, they rely on open table formats (OTFs). Unfortunately, they have many missing features inherent to their metadata designs, like no support for multi-table transactions and recovery in case of an abort in concurrent, multi-query workloads. This, in turn, can lead to non-repeatable reads, stale data, and high costs in productive cloud systems. In this work, we introduce LakeVilla, a complementary solution that introduces recovery, multi-query/table transactions, and transaction isolation to state-of-the-art OTFs like Apache Iceberg and Delta Lake tables. We investigate its transactional guarantees and show it has minimal impact on performance (2% YCSB writes, 2.5% TPC-DS reads) and provides concurrency control for multiple readers and writers for arbitrary long transactions in OTFs in a non-invasive way.
Analytics, Lakehouse, Object Store, Open Table Formats, Delta Lake, Iceberg, Concurrency, Transactions
copyright: acmcopyrightjournalyear: 2025doi: 10.1145/1122445.1122456conference: We do not know yet; …, 2025; Virtualbooktitle: Booktitleprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Lakehouses for analytical workloads. The past few years have seen the rise of a new type of data management system, called lakehouse (LH), which combines the benefits of low-cost, open-format data lakes and transactional data warehouses (Zaharia et al., 2021; Errami et al., 2023; Levandoski et al., 2024; Kraft et al., 2023; May et al., 2025). With the data gravity leading towards data lakes, LHs for analytical query processing became crucial for various tasks (Zaharia et al., 2021; Inc., 2023; Begoli et al., 2021; Hambardzumyan, 2023; Armenatzoglou et al., 2022; Services, 2020; Kuschewski et al., 2023; Levandoski et al., 2024; Schmidt et al., 2024; Haas and Leis, 2023). LHs use open table formats (OTFs) such as Delta Lake tables (Projects, 2024a), Apache Iceberg (Foundation, 2024a), or Apache Hudi (Foundation, 2021b) to define separate metadata layers on the object store for each table (Zaharia et al., 2021; Armbrust et al., 2020). Those formats implement transactions for single tables and operations, indexes, and time travel (Zaharia et al., 2021; Armbrust et al., 2020), but also lack support for critical features (Armbrust et al., 2020; Jain, 2024).
Missing Features in OTFs. This paper addresses the problem of how to support multi-query and multi-table transactions for LHs while providing strong ACID guarantees. More specifically, we provide native support for the following three missing features: First, the conflict resolution of OTFs cannot track concurrent changes, often resulting in multiple aborts and higher transaction costs. We refer to this feature as (i) Recovery. Second, OTFs today do not support cross-operation dependencies, thereby only allowing simple business logic of single query transactions(Zaharia et al., 2021; Armbrust et al., 2020; Projects, 2024b; Foundation, 2024g; Levandoski et al., 2024). We refer to it as (ii) Complex Transactions. Third, OTFs treat tables separately, leading to inconsistencies and dirty reads under concurrency, and we refer to it as (iii) Isolation. Unfortunately, the lack of support for these features prohibits concurrent writes in a system and can not guarantee properties like referential integrity across tables for concurrent reads. As a result, one cannot execute multi-query and multi-table transactions, forcing use cases for elastic compute or similar (Zaharia et al., 2010; Foundation, 2018; Sethi et al., 2019; Foundation, 2024h) to rely on synchronization methods that block the whole LH (Zaharia et al., 2021; Armbrust et al., 2020). For example, if we normalized our data into different tables like ’students’ and ’lectures’, some concurrent analytics could observe students taking an unknown lecture.
Table 1. Overview of existing formats and extensions and their supported features.
Multi-Query/Table: (i) – (iii)
System/Extension Single- query Txn Reco- very (i) Complex Txn (ii) Iso- lation (iii)
[…]
LakeVilla (our approach) $faThumbsOUp \faThumbsOUp \faThumbsOUp$
[…]
State of the Art. Despite multi-query and multi-table transaction support being a highly requested feature in the LH area (Jain and Databricks, 2024; Armbrust et al., 2020; Zaharia et al., 2021), prior work only partially addresses them. In Tab. 1, we summarized the state of the most notable approaches for multi-query and multi-table transaction support in LHs. In summary, none of them adapts to concurrent changes (cf. (i); Recovery) (Foundation, 2024e; Projects, 2024b; Jain et al., 2023; Foundation, 2024f) and some of them currently develop solutions for cross-operation dependency detection (ii) and global transaction isolation (iii). Despite known techniques like the SAGA pattern (Garcia-Molina and Salem, 1987) or MVCC (Bernstein and Goodman, 1983; Neumann et al., 2015; Bernstein et al., 2011; Liu et al., 2020; Iwabuchi et al., 2016) that ensure similar features in traditional database systems, all approaches listed in Tab. 1, add auxiliary components external to the object store but specific to a single OTF format (Jain and Databricks, 2024; licence, 2024; XTable, 2024). In contrast to the original idea of LHs as an open data warehouse (Errami et al., 2023; Zaharia et al., 2021), relying on these external services raises concerns for inter-operability (or unavailability) across vendors, potentially reenabling vendor-lock in. Hence, we aim to support multi-table and multi-query transactions without relying on any auxiliary structures.
LakeVilla (LV) is our extension that implements the three features for (i) recovery, (ii) complex transactions, and (iii) isolation to provide multi-table and multi-query transaction support under concurrency. LV defines a composable set of mechanisms that are non-invasive and compatible with existing OTFs. In contrast to prior work, LV’s mechanisms are integrated into the original protocol and enable each transaction to choose its desired guarantees individually. We ensure LV’s correctness by designing the techniques for the object store that guarantee all transactions result in a valid state of the raw OTF. This ensures a simple integration into existing LH systems and compatibility with clients running the raw OTF. More specifically, LV introduces three modular features: First, LV [Recovery] (LV [R]) uses logical snapshot version markers with transaction sublogs and physical and logical undo and redo information to define a greedy transaction model to support recovery (i). Second, LV [Complex Transactions] (LV [CT]) introduces a marker-based conflict and deadlock detection model among all tables to support complex transactions that include an arbitrary number of queries/operations (ii). Lastly, LV [Isolation] (LV [I]) introduces a global version log that checks global relations among tables and acts as a new layer on the object store above all tables (iii). We also demonstrate how each LV feature strengthens the ACID guarantees of LHs and discuss the performance implications.
Key insights. We show that our LV features guarantee serializability and linearizability for OTFs on the entire LH, not just individual tables or operations. The non-invasive integration of the LV [I] (reads) and LV [R, CT] (writes) can be done with minimal overhead of 2.5% (TPC-DS reads) and 2% (YCSB writes) and incur negligible and predictable costs for typical LH workloads, like analytics.
Our contributions: In this paper, we give an overview of the design area of open table formats in Sect. 2,
  • C1
    Define non-invasive features LV [R], LV [CT], and LV [I] for (i)–(iii) providing compatibility across all their combinations and the base OTF without vendor lock-in (Sect. 3),
  • C2
    Show that LV improves the ACID guarantees of OTFs involving multiple tables, queries, and concurrent writes especially guaranteeing Snapshot Isolation up to Serializability (Sect. 4),
  • C3
    Introduce a LV prototype with minimal overhead and predictable costs when executing our newly adapted YCSB-LH (Cooper et al., 2010), CAB-LH, (van Renen and Leis, 2023), TPC-DS, and microbenchmarks (Sect. 5).
Afterward, we discuss related work in Sect. 6 and conclude in Sect. 7.
2. The Lakehouse Design Area and Structure
Lakehouses (LHs) are an evolution from Data Lakes (Armbrust et al., 2020; Zaharia et al., 2021; Jain et al., 2023) for efficient processing of structured data in object stores using a query engine (Behm et al., 2022; Foundation, 2018, 2024h). LHs aim for comparably lower storage costs in data lakes while maintaining reasonable performance at the expense of data freshness for their expected workloads. The typical use cases of LHs are machine learning and analytical applications with concurrent writes from various data sources with unique characteristics (Zaharia et al., 2021; Armbrust et al., 2020; Foundation, 2024a; Projects, 2024a). This work focuses on analytical applications consisting of an initial load phase and mixed workloads of 40% continuous writes and 60% analytical reads. Writes are typically split between 80% inserts and 20% updates (Schmidt et al., 2024). Across SAP use cases from Ariba (SE, 2024c), S4/HANA (SE, 2024a), and SAP Sales Cloud (SE, 2024b), we found initial load sizes of multiple GBs up to x1 TBs. Workloads have ingestion rates of x103 kB to MB-sized events per hour and complex analytics with up to x102 to x103 concurrent users. The required data freshness lies between a few seconds to daily.
Open table formats (OTFs) and missing features. OTFs on the object store are the basis of LHs (Projects, 2024a; Foundation, 2024a, 2021b). Those formats add a separate metadata layer for each table, organize data, and control client concurrency. This enables advanced analytical features, easy adaption, and builds on open-source formats (Armbrust et al., 2020; Zaharia et al., 2021; Jain et al., 2023; Projects, 2024a; Foundation, 2024a, 2021b). Typically, OTFs use a file hierarchy and rely on atomic requests to the object store (Jain et al., 2023; Begoli et al., 2021; Zaharia et al., 2021; Chen et al., 2022; Armbrust et al., 2020). Two well-known OTFs (Zaharia et al., 2021) (i. e., Snowflake (Snowflake, 2024), Amazon EMR (Goswamia et al., 2022), BigQuery (Google, 2024; Levandoski et al., 2024)) are Delta Lake and Iceberg (Projects, 2024a; Foundation, 2024a). Delta Lake generates log files and checkpoints with unique ascending identifiers (Armbrust et al., 2020; Projects, 2024b). In contrast, Iceberg uses a tree structure to define a current table version represented by a unique branch (Suminto, 2023; Foundation, 2024g). However, those designs and other OTFs share the missing features acknowledged in prior work (e. g., (Zaharia et al., 2021; Armbrust et al., 2020; Jain and Databricks, 2024)) of which we cover the following:
  • (i)
    Recovery: Client crashes and aborts cause unreferenced files, additional costs, high latency, and redundant work.
  • (ii)
    Complex transactions: OTFs do not track dependencies across tables and operations, risking consistency and isolation for arbitrary long transactions.
  • (iii)
    Isolation: Reads might observe dirty data because each table handles commits and version assignments separately.
Different Solution Approaches. When addressing missing LH features, we observed four design directions described as questions:
(Q1) What concurrency mechanism should we use? Exclusive locking (Armbrust et al., 2020) prevents concurrent writes while aborting all conflicts. A branching solution, like Project Nessie (licence, 2024), isolates transactions in different branches (Foundation, 2024d). Lastly, presumably Delta Lake’s newly announced commit owner (Jain and Databricks, 2024) applies versioning-based conflict detection and resolution (Armbrust et al., 2020; Foundation, 2024a) known from DB research (Bernstein and Goodman, 1983; Bernstein et al., 1987; Neumann et al., 2015).
(Q2) Do we implement an internal or external solution? Many LH projects (XTable, 2024; licence, 2024; Jain and Databricks, 2024) add external services above the object store. This essentially provides a simple and performant solution. In contrast, a built-in solution within the OTF design without auxiliary architectures reduces dependencies on such services.
(Q3) Should we work with existing OTFs or redefine them? Delta Lake 4.0 update changes its original design and provides backward compatibility with its backfilling feature (Jain and Databricks, 2024). In contrast, adaptive approaches like XTable (XTable, 2024) and Project Nessie (licence, 2024) integrate into existing OTF architectures without noticeable structure changes, providing easier adaption on already deployed systems.
(Q4) For which type of transactions do we use which solution? Most LH projects are one-fit-all solutions affecting all transaction types (Armbrust et al., 2020; Jain and Databricks, 2024; XTable, 2024; licence, 2024). This contradicts a flexible solution, providing individual choice between guarantees. However, this limits its simplicity, as such solutions must offer modular building blocks.
3. LakeVilla
Our solution. We introduce LakeVilla (LV) as a versioning-based (Q1), built-in (Q2), adaptive (Q3), and flexible (Q4) OTF extension providing the missing features (i) – (iii). LV provides three modular features: LV [R] for (i) Recovery, LV [CT] for (ii) complex transactions, and LV [I] for (iii) isolation. Our design can seamlessly integrate into OTFs such as Apache Iceberg and Delta Lake without reformatting or reconfiguring the whole system.
3.1. LakeVilla [Recovery] (LV [R])
Why multi-query transactions abort. When operating with concurrent transactions, the key challenge to prevent aborts is adapting to changes in the table versions of the metadata layer. By default, OTFs apply optimistic concurrency control that checks for conflicts at the commit of a transaction (Projects, 2024b; Foundation, 2024f; Zaharia et al., 2021). This approach assumes that most operations do not conflict. However, OTFs also provide table-level granularity, meaning that all concurrent write operations on the same table will conflict without fine-grained conflict management. Hence, we propose LV [R] to provide recovery (i). This feature combines two techniques we designed for OTFs following our decisions for Q1–Q4: Our markers (technique #1) reserve snapshot IDs at a transaction’s beginning seamlessly and publish a transaction’s changes atomically. To adapt to concurrent changes and faults, we propose our sublog (technique #2) in LV [R].
(a) Apache Iceberg Integration
(b) Delta Lake Integration
Figure 1. Technique #1: Markers (green) reserve snapshot versions in the metadata layers
[…]
[Technique 1: We place file-based markers to reserve snapshot versions.]Technique 1: We place file-based markers in the metadata layer of OTFs to reserve snapshot versions. For Delta Lake, it appends to its log structure. The Iceberg integration places markers on a new path.
Technique #1: Markers We propose file-based markers within OTF metadata layers to prevent transaction starvation. For OTFs, they atomically reserve a table snapshot version during their transactions’ execution, thus blocking concurrent transactions from claiming the same version. We ensure their uniqueness and atomicity by using Put-If-Absent or conditional writes used in OTFs (Amazon Web Services, 2014; Armbrust et al., 2020). Additionally, we implement continuous checks for object stores without atomic requests accessing each marker’s HEAD to detect potential overwrites. If a client fails to claim a version or its marker was deleted by someone else, it aborts or claims a new marker. Markers are temporary objects that must be deleted or overwritten.
Marker integration. Their OTF integration shown in Fig. 1 is not arbitrary. In Iceberg, we place markers in the area of manifest lists and link to the snapshot. For Delta Tables, we append markers to the log directory. We show both approaches in Fig. 1(a) and Fig. 1(b) for marker ‘M1’, respectively. In both integrations, the marker becomes an exclusive lock (Bernstein et al., 2011; Liu et al., 2020; Iwabuchi et al., 2016; Bernstein et al., 1987) to prevent other concurrent transactions from claiming the same snapshot version (i+1).
Marker Compatibility. Following our design principle of flexibility (Q4), we propose markers should not enforce the same guarantees or new techniques to all transactions. Hence, our markers must fulfill their purpose while still being compatible with the original OTF protocols. LV’s markers use the extensible open-source formats (e. g., JSON) of the original logs/manifest and appear empty for the original protocols. At the same time, they are detectable for our marker-based clients with the data contained.
Marker Commit. Nevertheless, commits using markers should be atomically visible. So, we designed a commit phase that replaces these markers noticeable to legacy readers: Our client collects and summarizes all relevant information according to the OTF’s required instructions during a commit. If needed, it builds underlying structures, such as Iceberg’s manifest lists, and finally overwrites the placed marker with all the necessary instructions. This action effectively makes the whole transaction result of a table available even for the original OTF protocols. In case of an abort, our client overwrites a marker with a previous manifest list or an empty log entry to not disturb the original protocols.
(a) Physical
(b) Logical
Figure 2. Technique #2: Logical and physical Sublog instructions for ”UPDATE ¡Table ID¿ SET … WHERE key=¡key¿”
[…]
[Technique 2: Logical and physical SAGA instructions for ”UPDATE ¡Table ID¿ SET … WHERE key=¡key¿”.]Technique 2: Logical and Physical SAGA instructions for ”UPDATE ¡Table ID¿ SET … WHERE key=¡key¿”. The physical instructions explicitly command reading, adding, and removing a specific file. Logical operations describe the logic of an update by searching for a key and replacing the found file.
Technique #2: Sublog Structure. After introducing our marker system, we propose an in-flight transaction recovery system using our sublogs. This technique generates a structure of transaction logs for in-flight transactions utilized by Undos and Redos of a transaction. We designed our sublog structure and its sublogs from scratch because the underlying operations differ in file-based OTFs from database systems. Operations generate, swap out, and split files to create a complete version for the following snapshot. Delta Lake’s time-traveling (Armbrust et al., 2020) already defines basic redo information for OTFs in its logs. However, its definition is static, while a transaction must adapt in a conflict. Hence, we need a new logging type for redo operations in a concurrent LH environment. LV defines physical and logical sublogs to provide the necessary information for this technique. Figure 2 shows the generated LV sublogs in our physical and logical definition for the query ”UPDATE ¡Table ID¿ SET … WHERE key=¡key¿”.
Our Physical Sublogs describe the explicit changes made to the previous OTF table state, such as adding and removing files. We show an example of a physical sublog for the above query in Fig. 2(a). The first instruction in Fig. 2(a) reads the file with the specified key ¡key¿ for the selected table ¡Table ID¿, resulting in the physical path of the found data file ¡path_old¿ and the file’s contents. The respective compensation step repeats this operation, as the redo operation does not change the overall table state. For the next step, the client uploads the new data file under the newly generated path ¡path¿ and defines its removal for the compensation step. Lastly, the transaction removes the old data file at ¡path_old¿ and adds it for compensation.
Our Logical Sublogs encode the example update query as shown in Fig. 2(b). Its first sublog defines the search of the current file with the given ¡key¿ and stores the result in the client’s buffer. The respective compensation step logs the old value of the buffer “old_found” to restore the old buffer state. The second sublog references the current buffer value “found” and merges the referenced file with a new file at ¡path¿ to generate the updated version. For compensation, we replace the new file at ¡path¿ with the original file. A transaction must ensure that it executes a redo on all its logical sublogs before it commits.
Sublog: File-based Redos and Undos. An undo traverses the sublog structure from the newest to the oldest sublog, executing the compensation steps to restore previous states. For our example in Fig. 2, it effectively restores the transaction state before the update in both logging versions: For the physical instructions, it adds the old file at ¡path_old¿ and removes the updated file at ¡path¿ (compensation steps p+1&p+2). Furthermore, the undo mechanism can execute physical reads like (compensation step p). The logical version starts by replacing ¡path¿ with the path stored in its buffer (compensations step l+1). Internally, we implement replacements through the removal of ¡path¿ and the addition of the file at the buffered path (compared to physical instructions). Lastly, we must store the old buffer value ¡old_key¿ in the client’s buffer for the following undo instructions (compensations step l). In contrast to undo, redo operations traverse the sublog structure from the oldest to the newest file, executing the redo steps. This process effectively results in a client performing the same operations on the current table version.
LV [R]’s Interplay of physical and logical sublogs. The goal of our sublog technique (#2) is to be adaptive to changed environments due to concurrency. In LV [R], physical sublogs describe file operations already performed on the object store and are thus not adaptive anymore. Therefore, they have precedence over logical sublogs. LV [R] reserves physical sublogs for operations without preceding in-flight transactions. If we detect a preceding transaction by its markers, we allow switching from physical sublogs to logical sublogs. However, it is not allowed to switch back because this would destroy the precedence. Furthermore, logical sublogs try to greedily apply explicit changes as much as possible to improve performance. For example, our design sees blind appends to a table as non-conflicting. A client will generate the file and add a physical sublog, even in a logical structure. This mixture of techniques #1 and #2 enables higher throughput in the metadata layer and transforms our markers from resembling exclusive to shared locks.
LV [R]’s Conflict resolution and file-granularity. To determine whether to commit or abort, LV [R] uses the markers placed in the metadata layer of a table. The unique and distinct order among the atomic markers ensures an ordering among all operations within a table. For concurrent transactions, the conflict detection triggers as soon as a preceding transaction replaces its marker. The waiting client accesses the content of the new log and checks if the prior transaction changed the same files or keys using its logical file-based sublog. LV [R] solves such file or key conflicts by undoing the transaction to the first operation where a conflict was detected and triggering a redo on the new table state. However, logical conflicts like dependencies across transactions or tables might trigger more prolonged recovery procedures, potentially completely re-executing the transaction. If an error appears multiple times, we added a retry counter that interrupts recovery if necessary and aborts the transaction. As OTF logs and snapshots operate on file references, this combination gives us file granularity for all transactions using LV[R].
LV [R]’s Client failure recovery. Besides our conflict resolution during commit, our LV [R] guarantees recovery from other causes of failure. Some common issues for OTFs revolve around upload/download failure, client crashes, or similar problems. In case of such failures, they might leave the LH metadata layer in an inconsistent or invalid state. Further, adding or removing files for updates or using markers and other LV techniques amplifies this problem. Hence, LV [R]’s sublogs are also a safety mechanism that allows clients to undo all failed transaction changes even if they did not initiate them. To guarantee recovery in such situations, LV[R] requires that sublogs can only removed after clients replace their marker with a snapshot/log. Currently, LV uses timeouts and counters for failed transactions to detect crashed clients with active markers. If a client decides to free the markers of a potentially crashed client, it does not delete the sublog. This rule enables the crashed client to automatically claim a new marker and continue the transaction.
3.2. LakeVilla [Complex Transactions] (LV [CT])
Multi-Query and Multi-Table Transactions in Lakehouses. Building upon LV [R], we implement multi-query and multi-table transactions for LHs. Re-applying the marker idea only partially implements complex transactions as multi-table transactions break global dependencies due to the independent metadata layers for all OTF tables. The metadata layer of one table would not be able to recognize any dependencies between tables, thereby enabling inconsistencies, deadlocks, and partial aborts. We avoid such problems in LV [CT] by combining our markers (technique #1) with a conflict detection model (technique #3): We demand a transaction must announce its global presence and check table compatibility using its markers.
Figure 3. Technique #3: A marker-based conflict and deadlock detection model. This figure shows a conflict between tables 1 and 2 and how T1 recovers from it.
[…]
[Technique 3: A marker-based conflict and deadlock detection model. This figure shows a conflict between tables 1 and 2 and how T1 recovers from it.]Technique 3: A marker-based conflict and deadlock detection model. This figure shows a conflict between Tables 1 and 2. T1 detects a conflict by reading T2’s markers. It frees its marker and places a new one to resolve the conflict.
Technique #3: Marker-based Conflict and Deadlock Detection. While a table-level lock and redo and undo information can prevent conflicts and dependencies within a table (LV [R]), we still can not capture dependencies between tables. For multiple tables, we can encounter a deadlock or cause inconsistencies when placing more markers into these tables and waiting from preceding transactions (compare deadlocks for locks). We track dependencies by introducing our deadlock and conflict detection for our OTF markers (technique #1) and define a global conflict and deadlock detection model. Our goal is to track dependencies across tables without causing deadlocks among transactions.
Detection Strategy. This technique requires transactions to acquire markers in every table accessed. Markers placed in tables must contain information about the respective transaction’s dependencies and concurrently check for changes in detected markers of other transactions. We explain our strategy with the example in Fig. 3: Transactions T1 and T2 start working on different tables and place their markers in the respective metadata layer. These transactions can be considered unrelated at this point. Later, T2 accesses Table 1 and places its marker behind T1’s marker. When initializing the new marker and reading the table state, T2 recognizes T1’s marker. As this relation does not violate its observed dependencies, T2 doesn’t abort. It updates all its lists of Tables 1 and 2 by overwriting each marker. T2 depends now on T1. Afterward, T1 accesses table 2, places its marker in the metadata layer of table 2, and checks for predecessors. T1 detects T2’s marker and retrieves its data. It compares the opened tables with T2’s list and checks for potential deadlocks and conflicts. Because T2 also lists table 1 as accessed and T1 is not aware of the ordering T1 depends on T2, it concludes that a deadlock might occur and triggers a marker shift. Hence, T1 frees its old marker in Table 1 by overwriting it with an empty manifest/log and places a new marker ”Marker T1’” into the respective metadata layer. After confirming that the dependencies match (also possibly readjusting some changes in Table 1), T1 continues its operation. If T2 decides to access another table and finds a conflict contradicting its internal observations, it will reevaluate this observation. Here, it checks the origin of this observation (here: Table 1) if the situation changes. Depending on the result, it updates its internal observations or solves the conflict.
LV [CT]’s Integration For multi-query transactions on the same table, we revisit the marker idea (technique #1) and extend it with technique #3. Like before, LV holds a marker to lock a table until the entire multi-query transaction is processed and overwrites with a real manifest/log to unlock (compare locks (Bernstein et al., 1987)). LV [CT] exposes information by allowing an in-flight transaction to claim access to new tables with multiple markers and using marker-based conflict detection. If a client detects a deadlock, it resolves the issues with technique #3’s marker shift. Furthermore, when successful in all tables, a commit triggers all table-level commits required for a transaction by replacing all markers and commits.
LV [CT] concurrency and delay handling. Due to the overhead of concurrent marker checks, marker generation, and general network delays, LV [CT] can not guarantee that all markers are up to date at any time. Hence, our file-based deadlock detection model must consider such uncertainties. We introduce a best-effort strategy: the concurrent checks to dependent markers searches for changes in dependencies. However, a transaction might be unable to detect the complete dependency cycle due to client crashes or faults. If a transaction is stuck waiting for other transactions, it triggers a marker shift after a timeout. After a timeout triggers a marker shift, the transaction aborts if a second timeout occurs. This strategy ensures that transactions do not hold their markers indefinitely, even under uncertainty, blocking the entire system.
Proof Sketch: Conflict Resolvement We demonstrate that LV [CT], combining techniques #1 and #3, effectively exposes and resolves conflicts across tables. From LV[R], we know that markers represent a running transaction within a table. They are unique for each transaction. As an LV [CT] client places a distinctive marker in each accessed table and waits for preceding ones, deadlocks can only occur if concurrent clients form a circular wait scenario across tables. Hence, if a transaction is part of such a cycle, it atomically places a marker behind another transaction’s marker, which waits for a different marker in another table. LV [CT] exposes its dependency list within all its markers, informing waiting transactions. At the same time, a client checks all markers it is waiting for to construct a dependency graph. If some given list contains the transaction’s ID or the provided information forms a cycle, an LV [CT] client is part of the cycle and triggers a marker shift. In case of larger cycles that might not be visible in the provided data, we use a timeout while waiting for markers to guarantee marker shifts breaking such cycles. Hence, we created a defensive approach that triggers shifts and potentially redos over too much wait. Performing a marker shift (technique #3) effectively frees all markers within such a cycle and claims newer ones. Hence, the marker shift inverts the relation between the transaction performing the shift and all waiting transactions. Therefore, during their concurrent checks, waiting transactions will realize that the marker is freed and proceed, breaking the cycle.
3.3. LakeVilla [Isolation] (LV [I])
The need for a global Lakehouse Layer. We introduce LV [I], defining a global layer that manages commits and global dependencies (technique #4). This change guarantees globally consistent versions at all times and global isolation (iii). Furthermore, it expands steps from the original OTFs and MVCC, like version assignment and validation (Bernstein and Goodman, 1983). Our approach minimizes cost and avoids slow requests like LIST to the object store.
Figure 4. Technique #4: A global version log tracking all valid snapshots. T2 succeeds and updates snapshots of Tables 3–5.
[…]
[Technique 4: A global version log tracking all valid snapshots. T2 succeeds and updates the snapshots of Tables 3, 4, and 5.]Technique 4: A global version log tracking all valid snapshots. The conflicting transactions T1 and T2 both enter the file-based validation phase. While T1 fails and is considered incompatible, T2 succeeds and updates the snapshots of Tables 3, 4, and 5.
Technique #4: Global Versioning After introducing our techniques to reserve snapshots, recover, and detect global deadlocks, we propose a global layer for object stores that manages all table snapshot versions to track and isolate transactions globally. Our global version log summarizes all tables’ most recent valid versions. If a client accesses a table, it is bound to the versions received by the global version log and can only use table versions equal to or smaller than the specified version. The example in Fig. 4 shows a global version log that tracks four tables and their newest snapshot versions. T1 and T2 both start their execution and retrieve the version combination (Table 1, Ver. 5), (Table 2, Ver. 3), (Table 3, Ver. 2), and (Table 4, Ver. 6) from the global version log.
Version Validation on the object store. To check the transaction’s compatibility with the current global version, we propose a file-based validation phase for the commit in OTFs, with a similar purpose as MVCC (Bernstein and Goodman, 1983; Neumann et al., 2015). This validation of a transaction starts by checking and atomically copying its version changes to a shared key on the object store. If the transaction succeeds, it iterates over its transaction version log and compares the version numbers of the current global version log. However, table versions in LHs are different. We can distinguish between a table’s read version (snapshot version we got from the first access of a table), hidden version (all versions created after a client received its read version), write version (id of the created log/marker), and log version (table version during validation). Using those abstractions, we define conflict rules for our global version log. We consider a transaction as compatible if it does not fulfill any of the following rules:
  • R1
    read version ¡¡ hidden version (outdated data)
  • R2
    write version ¡= hidden version (log number already taken)
  • R3
    log version + written logs != write version - written logs (lost update)
In summary, all of our rules have different purposes: R1 ensures that a read query does not access data that is too outdated from the current state. Further, R2 ensures that no log version is assigned twice, handles concurrency, and commits during transaction execution. Lastly, R3 guarantees the isolation with other transactions by only committing its logs. If a transaction is compatible, it executes the table-level commit of each accessed table before atomically announcing the new snapshots. In LV [I]’s snapshot isolation mode, only R2 and R3 are enforced, making LV [I] only check write sets.
LV [I]’s Read sets. We define our version log by using the well-proven read-and-write set abstraction. However, in LHs, we deal with table and file versions that must follow rules R1 – R3. For read transactions to retrieve globally consistent versions, they read the global version log and skip the snapshot assignment from their original OTF protocol. This change effectively reduces the number of requests to the metadata layers of single tables and replaces expensive requests (e. g., LIST for Delta Lake (Armbrust et al., 2020)). LV [I]’s read sets contain all tables read by a transaction. Hence, the LV [CT] transaction does not update such table versions. During its execution, we implement checking for concurrent changes in each table and updating hidden versions. In the serializable protocol, reads must pass the validation phase. Especially R1 was designed to uphold a configurable distance between the read and hidden versions. However, LV [I] can be relaxed to gain write serializability by automatically letting each read-only transaction pass validation and only incorporating the hidden version if necessary.
LV [I]’s Write sets. An LV [CT] write transaction follows the proposed strategy of technique #4. It retrieves a globally consistent read version by accessing the global version log and executes all its operations independently. Similar to reads, writes check for the hidden version of a table. A transaction’s write set contains all tables where it effectively created at least one temporary log entry. The respective validation checks its changes and atomically updates the global version log. To guarantee consistency, the transactions targeted write version number must be larger than their current log version, and the difference must match the number of added log files. LV [I] defines two special cases for writes in its validation: Updates and blind appends. Updates have to be read and written on the same table. Hence, they will have the same table in their read-and-write set. LV [I] enforces R1 – R3 on those sets independently and concurrently, making them the most work-intensive operations. On the other hand, blind appends only insert new data and are not bound by any restrictions. Hence, their order of operations is irrelevant. LV [I] implements a special flag for such operations that allows the transaction to take the following valid log ID directly in case of a validation failure.
Constructing an LV [I] version log. When accessing an OTF environment with LV [I] for the first time, the global version log does not exist. Hence, we construct this layer on demand using a mixture of the original OTF and technique #4: If an LV [I] client recognizes that no global version log exists, it proceeds by individually deducing its version of each accessed table. When committing, the transaction atomically enters the validation phase, accesses the empty global version log, and generates all entries based on its accessed table version. Other transactions accessing tables not contained in the version log will perform a similar mixture of claiming the version of the respective table themselves and adding the table during their validation. LV [I] defines adding a previously unknown table as always compatible because it can not check R1 – R3. When executed for a while, this procedure generates the global version log without disrupting other clients and replaces the version assignment proposed by technique #4.
LV [I]’s compatibility to other levels and original protocols. When using LV [I] alongside other client configurations, we face the problem of them ignoring the global version log. We ensure the compatibility of LV [I] by defining different behaviors for read and write transactions: Reads can still operate as described for technique #4. Such transactions do not conflict with concurrent transactions as they do not advance a table’s version number. In contrast, write transactions must be aware of such changes as concurrent non-LV [I] writes advance the version number without notifying the version log. If an LV [I] transaction decides to write to a table, it will check if the next possible version is available. If it detects a snapshot/log/marker of the same version, an inconsistency to the global version log occurs. It proceeds by reading the previously unknown logs, setting its hidden version to the newly read ID, and writing its results using the hidden version. During its commit, it counts the unknown log entries as it changes and commits them to the global version log if its initially assigned read version still equals the global version log. Otherwise, another concurrent transaction has already been committed using this version, and we must abort. Finally, to make writes using LV [I] visible to all clients, LV [I] still generates a valid log entry for each table.
4. LakeVilla’s Guarantees
After introducing the design of all LV features, we investigate their ACID guarantees summarized in Tab. 2. This section introduces the informal intuition behind individual proofs.
Table 2. Highest ACID guarantees achievable for OTFs + LV
Iceberg/ Delta Lake LakeVilla [Recovery] LakeVilla [Complex Txn.] LakeVilla [Isolation]
A OTF structure (ST) Atomic Markers (ST) Atomic Markers (ST) Version Validation
C Eventual (SQ,SW) Causal (ST) Causal Lineariz- ability
I Serializable (SQ,ST) (Databricks, 2024; Jain et al., 2023) Serializable (ST) Repeatable reads Serializable (LV)
D Object store guarantee Object store guarantee Object store guarantee Object store guarantee
Single Query only: (SQ)  Single Table only: (ST)
Single Writer only: (SW)  LakeVilla only: (LV)
Atomicity of LakeVilla. All LV Features guarantee table-level atomic and durable transactions as shown in Tab. 2. Atomicity revolves around the atomic markers in LV [R] and LV [CT]. Given the metadata structure of an OTF table, a transaction overwrites its marker with all changes made during its commit. Hence, any concurrent transaction will see either the ”empty” marker (see Sect. 3.1) or the complete log file. Because a marker is a single file, all changes of a transaction become visible at the same time when overwritten. However, this is limited to single tables due to the district metadata layers of OTF tables. In contrast, LV [I] guarantees Atomicity through its validation step, which uses atomic operations to allow only one client to modify the global version log. The head of the global version log is a single file that can be overwritten by a validated transaction in one operation, globally exposing all its changes simultaneously.
Durability of LakeVilla. Similar to the original OTF protocols, the guarantees of object stores and a client successfully uploading all new files before generating the log file (e. g., compare (Armbrust et al., 2020)) achieve Durability for LV [R] and LV [CT]. Further, a transaction only updates LV [I]’s global version log if it successfully enters validation and all tables’ metadata and data files are written. Hence, LV [I] uploads all data files to the object store before committing and can rely on the object store guarantees to provide Durability.
4.1. Consistency
The Database Consistency of LakeVilla depends on the base OTF. Due to their metadata structure, Delta Lake and Iceberg offer ACID transactions for single tables (Projects, 2024b; Foundation, 2024a; Armbrust et al., 2020). All OTF transactions are ordered by the table’s ”chain” or ”tree” of the respective OTF metadata layer. Due to the atomic operations used (PUT-IF-ABSENT, Conditional Writes, extra service), each transaction claims a distinct spot in the metadata layer - defining the order of all successful transactions within the table. Because LV’s features keep this design and always lead to a valid state of the base OTF, all its features guarantee table-level consistency. Since OTFs are used in distributed settings and to clarify the feature’s consistency guarantees, we investigate the distributed consistency levels of each feature.
LV [R]: Causal Consistency (Single-Table). The causal consistency of LV [R] among clients using this feature is implied by its atomic markers. These markers reserve a snapshot version in the respective metadata layer and define an order between transactions. The operations of a transaction are already ordered. As clients writing with the original protocol might ignore markers, LV [R] checks for any logs written after its maker and aborts. However, LV [R] clients are not linearizable because write operations on different tables have no order among them.
LV [CT]: Causal Consistency. LV [CT] provides causal consistency globally. A client’s operations are ordered by their execution order (related to the fact that they originate from the same client). Concurrent operations accessing the same table(s) are also related because they access the same data items (=tables). We define the order of such operations within a table using atomic markers. If a concurrent operation holds an earlier marker, all its operations are ordered beforehand; otherwise, all its operations are ordered after. However, LV [CT] is not serializable because it does not define an order among write operations on distinct table sets.
LV [I]: Linearizability. Regarding LV [I]’s isolation level, we designed the validation process to be atomic. Hence, all write transactions must pass the validation sequentially, defining a global order. However, overlapping read-only transactions that retrieve their versions from the global version log during an active validation phase will work with outdated data. When they enter validation, they will likely succeed depending on the interval chosen for LV [I]’s R1. Hence, the read-only transaction must be ordered before the transaction processing the validation previously. This guarantees us a serial global history, not the order transactions committed. Hence, LV [I] guarantees Linearizability.
4.2. Isolation
Base OTFs: Serializability (Single-Query, Single-Table). OTF formats guarantee serializability limited to single-query and single-table transactions. We observe that all OTF transactions define a sequential order among their operations. Further, OTFs like Delta Lake define a serial order between all log files in their metadata layer (Armbrust et al., 2020). Since OTFs only feature single-query and single-table transactions, all transaction operations are contained within a single log file, providing serializability or snapshot isolation (Databricks, 2024; Jain et al., 2023). As a result, the execution order of all transactions equals a serial execution. Their operations can be ordered transitively to form a serial history between all operations on the same table.
LV [R]: Serializability for single tables. LV [R] extends the guarantee of the base OTFs by adding atomic markers placed during a transaction’s start. Hence, all distinct transactions operating on the same table initially get a unique version. Using atomic object store operations guarantees uniqueness among the assigned versions for each transaction. This defines the version a transaction will use during its execution. Further, LV [R]’s marker technique guarantees that all transaction operations are contained in the same log file of the reserved version. This effectively defines a sequential ordering of transactions and their operations on the same table. Like in the base OTFs, transactions are sequentially ordered by the OTF structure and operations within their transaction. Hence, because the markers ensure correct positioning of all operations within the metadata layer, they uphold the structure of the base OTFs. At the same time, markers make it possible to put any number of operations into the correct position, dropping the original OTF’s limitation on single-query transactions.
LV [R, CT]: Serializability. With LV [R]’s shift to a table-level focus, LV [CT] guarantees repeatable reads of the entire LH. Like LV [R], LV [CT] places atomic markers in a table to claim a version for a transaction. However, it extends this model to claim individual table versions in all accessed tables. The defined deadlock detection model and marker shifts guarantee that the same order among related transactions is upheld across tables. A client accessing a new table will detect any previous markers within. This defines the order in which the transaction depends on all other transactions of the discovered markers. If these dependencies form a cycle or timeout, LV [CT]’s marker shift inverts the previous dependency in all other tables, ensuring it is upheld. If the newly found dependencies do not violate previously found ones, no marker shift is required, and local dependencies become globally enforced on all tables. Hence, because a client must redo some of its operations after a marker shift, LV [R, CT] guarantees serializability globally.
LV [CT]: Repeatable Reads. However, without LV [R], LV [CT] enables phantom reads after a marker shift. A transaction executing a query before a marker shift may receive a different result set for ranges afterward. Caused by the marker shift, the table versions between transactions were reversed, and newer changes that might have affected the result set were adapted. However, by remembering previous reads before the marker shift, reusing the old read versions, and checking for compatibility during commit, LV [CT] can prevent non-repeatable reads. Hence, LV [CT] guarantees the repeatable read isolation level.
LV [I]: Serializability (LV [I] only). LV [I] provides the isolation level of serializability. Its global version log defines a total order between all committed transactions. Each transaction must enter its atomic validation phase and check its changes using R1 – R3. It can only succeed if no other committed transaction changed its write set; otherwise, it must abort. Hence, the order of successfully validated transactions gives a sequential global order. Due to the atomic design of the validation phase, LV [I] guarantees that previously validated operations will be ordered before and after the current transaction. Similarly, as the global version log is a global layer for the entire LH, it defines relations between all transactions, giving Serializability. Hence, LV [I]’s ordering is not limited to table level. If transactions only check R2 – R3 during validation, we get Snapshot Isolation as we only check write sets. However, concurrent operations not using LV [I] ignore the global layer. To avoid inconsistencies, LV [I] uses the hidden version to adapt to such changes and still perform its validation, guaranteeing Serializability/Snapshot Isolation among LV [I] transactions.
4.3. Experimental Validation
Table 3. Shortened Hermitage Results (multi-query) for Spark using Delta Lake (Spark [DL]) or Iceberg (Spark [IB]) and LakeVilla [LV]. Anomalies described with ”-MT” are multi-table versions of the original scenario. We tested all anomalies and show the ones with the biggest differences here.
Anomaly Spark [DL] Spark [IB] LV [R] LV [R,CT] LV [R,CT,I]
G0-MT (Write Cycles) $faThumbsOUp * \faThumbsODown \faThumbsOUp \faThumbsOUp$
G1a (Aborted Reads) $faThumbsODown \faThumbsOUp \faThumbsOUp \faThumbsOUp$
G1b (Interm. Reads) $faThumbsODown \faThumbsOUp \faThumbsOUp \faThumbsOUp$
G1c-MT (Circ. information flow) $faThumbsODown * \faThumbsODown \faThumbsOUp \faThumbsOUp$
OTV (Obs. Txn. Vanishes) $faThumbsODown \faThumbsOUp \faThumbsOUp \faThumbsOUp$
P4 (Lost Update) $faThumbsODown \faThumbsOUp \faThumbsOUp \faThumbsOUp$
Concurrent Reads (Linearizability) - - $faThumbsODown \faThumbsOUp$
[…]
*=implied by multi-query results and limitations
To provide further validation of LV’s guarantees, we experimentally check Spark and our prototype using LV [R], LV [R, CT], and LV [R, CT, I] for the anomalies described by Hermitage (Kleppmann, 2014). A selection of the results is shown in Tab. 3 with additional multi-table versions of the same anomaly (marked with -MT). The original Delta Lake (Spark [DL]) and Iceberg (Spark [IB]) were included despite their restriction on single-query transactions, which led them to fail most multi-query anomalies. An exception is G0-MT (write cycles), avoided by their limitations and thus marked uniquely (G0-MT). In contrast, LV [R] prevents table-level anomalies like lost updates (P4), aborted reads (G1a), or Intermediate Reads (G1b) and adapts to aborted/crashed transactions (OTV). When adding LV [CT], the design can resolve conflicts across tables like Write Cycles (G0-MT) and Circular information flow (G1c-MT). Further, LV [I]’s global isolation guarantees concurrent transactions to observe globally consistent versions across tables. In summary, these results confirm our targeted guarantees.
5. Evaluation
Overview. We start by analyzing the different phases of the LV Features and comparing them to Spark’s Delta Lake (Foundation, 2018; Projects, 2024a). We investigate their effects under concurrent workloads using YCSB [^1] YCSB-LH available: https://github.com/goetztj/YCSB-LH (Cooper et al., 2010; Ren et al., 2024) and test different feature combinations based on the cloud analytics benchmark (CAB) (van Renen and Leis, 2023). We conclude by running LV on selected analytical queries from TPC-DS and estimating the added overhead for systems like Spark.
5.1. Experimental Setup
Object Store: MinIO. For all experiments, we use MinIO (Foundation, 2024b) on a cloud instance with an Intel(R) Xeon(R) Platinum 8260 CPU with 48 cores at 2.40GHz with 524GB of RAM and 5TB of SSD storage running Linux. MinIO provides strict consistency (Sarrel, 2024), conditional writes (Amazon Web Services, 2014), and analytical features for the experiments. It is accessible through the AWS S3 interface and behaves like AWS’s offering.
LakeVilla Prototype. Our C++ LV prototype [^2] LakeVilla prototype available: https://github.com/goetztj/LakeVilla. implements a baseline Delta Lake client (LV [DL]) with all modular features (LV [R], LV [CT], LV [I]). It runs on an Intel(R) Xeon(R) Platinum 8260 CPU at 2.40GHz with two sockets of 64 cores and 1048GB RAM. We chose eight threads per LV client unless stated otherwise.
Spark: Delta Lake. Spark 3.5.3 runs with Hadoop 2.12, Delta Lake 3.0.0 (checkpoint interval 10), and Hadoop AWS 3.3.4. We minimized caching to observe the OTFs’ effects on the system. Our HIVE 3.0.0 catalog and its PostgreSQL backend run in two separate docker containers. Spark uses Java 11 and runs distributed on 9 machines with the same setup as above. The coordinator runs on one server with specifications like our MinIO server and shares its resources with the catalog store and benchmark drivers to reduce the networking effects between these components. For workers, we configured up to two instances on eight servers using an Intel(R) Xeon(R) Platinum 8260 CPU at 2.40GHz with 8 cores each. Each worker was assigned 64GB RAM and eight cores.
5.2. Phases of OTFs
(a) LakeVilla Prototype
(b) Spark
Figure 5. Phases for read and write workloads of a single transaction for YCSB queries.
[…]
[Phases for read and write workloads of a single transaction for YCSB queries.]Phases for read and write workloads of a single transaction for YCSB queries. The prototype’s reads are comparable to Spark, while Spark’s writes spend more time on the OTF version assignment.
Comparison of Spark and LakeVilla. We implemented our LV prototype as a separate C++ client. Thus, we first compare the differences between it and other systems. For this purpose, we ran YCSB sequentially with Spark using Delta Lake (Spark [DL]) and LV [DL]. If not stated otherwise, we use the default YCSB of 1000 operations on 1000 entries with row sizes of 1KB (Cooper et al., 2010). We logged the different phases of both implementations and calculated both the average latency for reads and writes, and how much each phase contributes to the overall latency. We expect both systems to display similar behavior. We show the results in Fig. 5. Spark [DL] and LV [DL] indeed behave similarly for reads. They spend most of their time accessing and reading the metadata layer of the YCSB table (95% and 93%, respectively) and the rest to load the requested data.
We observe a notable difference for the write requests. Spark [DL] ’s write latency is dominated by accessing the metadata and generating a new log entry within the metadata layer (69% and 28%). The actual operation of uploading new data is just a fraction of its latency. In contrast, LV [DL] ’s numbers indicate more time spent on data generation, while access to the metadata layer and log generation play a minor role. This difference is critical to interpreting the subsequent results in our evaluation so we can reliably estimate the impact of integrating LV in systems like Spark.
The different LV Features. Next, we evaluate the impact of each LV feature on latency. We use the same setup as before and show our results in Fig. 5(a). First, we analyze the overhead of placing LV [R] ’s and LV [CT] ’s markers (technique #1). We hypothesize that it affects the latency, especially for reads. In Fig. 5(a), writes show minimal impact as they align with metadata layer operations in the original OTF (LV [DL]). Further, freeing those markers is fused with the write transaction’s log placement, compensating its effect as seen in Fig. 5(a). In contrast, reads experience a 20% latency increase due to added OTF synchronization, with an additional impact when we free the markers post-read. This indicates that LV [R] and LV [CT] are better suited for write operations.
Next, we analyze the impact of LV [R] ’s sublog (technique #2) and LV [CT] ’s deadlock detection (technique #3) on latency using the same setup. Because both techniques are expected to work concurrently alongside the standard operations, we expect minimal impact on latency, which is also confirmed by Fig. 5(a). The time spent by the prototype on the actual operation in LV [R] and LV [CT] matches the baseline (LV [DL]). LV [R] ’s sublog generation is distinct and does not impact the baseline operation. Further, LV [CT] ’s deadlock detection does not detect problems as YCSB uses only one table. Hence, under non-concurrent workloads, neither technique impacts read or write latency.
Lastly, we expect LV [I] ’s global version log (technique #4) to have a noticeable but smaller impact on latency in the single-table YCSB setup. As seen in Fig. 5(a), it replaces the base table’s local version assignment, adding a 4ms overhead for reads and writes compared to the baseline. Further, even though the validation phase takes longer for writes, it more significantly affects the latency of the reads. In LV [I] ’s WriteSerializable mode, version assignment is skipped for reads, thereby shortening their latency. In summary, the global version log introduces overhead for all operations in a sequential, single-table setting. However, the WriteSerializable mode of LV [I] can mitigate this impact in such scenarios.
Figure 6. YCSB when using Spark (Delta Lake) and LV [R].
[…]
[YCSB when using Spark (Delta Lake) and LV-R.]YCSB when using Spark (Delta Lake) and LV [R]. Spark shows small throughput improvements and more aborts for more concurrent clients. LV [R] committing after each operation shows a slightly higher throughput than Spark with no aborts and increases with more concurrent clients. When bundling all operations in a single transaction, LV [DL] shows a high throughput with no aborts but a decrease in throughput for more concurrent clients due to the required redos.
5.3. Concurrent Workloads
OTFs and Concurrent Workloads. To evaluate the effects of concurrency and supporting multiple writers, we execute YCSB (Ren et al., 2024; Cooper et al., 2010) with 1, 8, and 64 concurrent clients and compare Spark [DL] with LV [R]. Spark was allowed up to 10 retries for a failed query. We evaluated LV [R] using single-query transactions (1Q Txn) and multi-query transactions that bundles all operations of a client (single Txn). We expect the throughput to increase for all workloads and systems, but also that the frequency of the conflicts will impact the behavior both through aborts by affecting the throughput. Figure 6 shows our results for YCSB A, B, C, D, and F. YCSB E was not tested because our current LV prototype does not support scans. Confirming our hypothesis, Spark experienced a slight increase in throughput, but with abort rates rising alongside the client count. Surprisingly, we also measured a small number of aborts for the read-only workload YCSB C, which we found to be an internal problem in Spark. In contrast, LV [R] had no aborts and outperformed Spark in throughput, particularly when bundling operations, which reduces version assignments per client. When we include the commit latency of LV [R] (single Txn) in our throughput calculations, we can also measure the impact for each workload type: For example, write-heavy workloads (A, F) experience a drop in throughput when we increase concurrency due to redo operations triggered at commit time. In contrast, executing LV [R] (1Q Txn) degenerates into normal Delta Lake with built-in retries, resulting in a throughput closer to Spark. As to be expected, bundling multiple operations into a transaction is advantageous for throughput as it minimizes the overhead of the version assignment phase. However, when overused, it may lead to penalties due to redo operations. In conclusion, we show that concurrency for writes is possible with LV [I], but users must be mindful of the suitable transaction sizes. The more writes a workload has, the less requests should be bundled into single transactions.
(a) 50% reads
(b) 95% reads
(c) 100% reads
Figure 7. YCSB (8 clients): Comparison of LV [I] and LV [R]
[…]
[YCSB (8 clients): Comparison of LV-I and LV-R]YCSB (8 clients): Comparison of LV [I] and LV [R]. LV [R] is faster for YCSB’s small reads and writes but more unstable than LV [I].
LV[R] and LV[I]. Lastly, we use YCSB with varying workload types to compare the impact of LV[R] ’s markers and LV [I] ’s global version log on latency under concurrency. Again, we bundled all operations into one transaction per client and ran YCSB with eight clients. LV [I] ’s hidden version interval was set to two. We hypothesize that LV [R] will be more suitable for mixed workloads, whereas LV [I] will be better for read-heavy workloads. In Fig. 7, we report this experiment’s read and write latencies. They confirm our hypothesis in the following way: LV [R] provides better latency at the cost of more outliers. These are caused by marker acquisition failures, clean-up, and transaction retries. This is independent of the workload type, as reads and writes synchronize in the metadata layer. Conversely, LV [I] adds higher overhead overall but stabilizes the read operations, which can now only conflict with extremely outdated table versions (Fig. 7(b) and Fig. 7(c)). In write-heavy workloads (Fig. 7(a)), LV [I] caused more outliers due to failed validations requiring full transaction retries. In summary, LV [R] optimizes concurrent transaction synchronization for OTFs with faster performance but higher variability, while LV [I] stabilizes the reads under concurrency without introducing writer’s congestion.
5.4. Benefits and Trade-offs for Transactions
Figure 8. Table Scaling of LV [CT]. The client in the table scaling experiment wrote one row to every opened table.
[…]
[Table Scaling of LV-CT. The client in the table scaling experiment wrote one row to every opened table.]Table Scaling of LV [CT]. The client in the table scaling experiment wrote one row to every opened table. LV [CT]’s latency requires at least 2 threads per table to compensate for its overhead. Afterwards, we see a linear increase.
LV [CT] ’s trade-off. Next, we evaluate the strengths and performance trade-offs of LV [CT], which supports multiple tables. Hence, unlike LV [R], its overhead depends on the number of accessed tables per transaction. We scaled the number of tables, following the customer table definition of TPC-H. All transactions inserted a single row into each table to reduce operation overhead but ensure it triggers the LV [CT] ’s conflict detection across all tables. The LV clients use 64 threads, and our hypothesis is that a transaction’s commit latency will increase with the number of tables due to more metadata layers and markers being involved.
Our results can be seen in Fig. 8. As expected, we can observe the overhead of LV [CT] ’s conflict detection, once we access more than one table. However, the transaction latency stays stable (at ca. 1s) until we run transactions accessing more than 32 tables. The parallelization in our LV prototype (and assigning 64 threads to our engine) causes this effect. Our prototype uses two threads per table for LV [CT]: one to execute the operations and the other to track dependencies to other tables. Hence, as soon as we run out of threads, some threads will need to manage two or more tables simultaneously, causing increased latency.
Additionally, also the commit latency increases with each accessed table. The tracking of dependencies is the overhead of LV [CT], but this does not impact the overall latency for transactions with less than 32 tables. The expanded commit procedure of generating the log/snapshot entries and overwriting the markers for each accessed table separately only impacts this feature minimally.
(a) Sequential Runs
(b) Concurrent reads and writes
Figure 9. CAB-LV: Level Switching of read and write streams with their respective throughput on TPC-H data.
[…]
[CAB-LV: Level Switching of read and write streams with their respective throughput on TPC-H data.]CAB-LV: Level Switching of read and write streams with their respective throughput on TPC-H data. The best combinations for reads are LV [DL] and LV [I], while large writes benefit from LV [R] and LV [R, CT]. These recommendations hold under concurrency.
CAB-LV: Feature Combinations We now proceed with more complex workloads to investigate the effects on performance when we switch from one LV feature combination to another. To assess that, we adapted the Cloud Analytics Benchmark (CAB) (van Renen and Leis, 2023) for LHs and distinguish between read and write streams. Read streams execute the streams generated by the original CAB generator consisting of TPC-H queries. The write streams either insert up to 100 rows to a random table per query (80%) or issue update requests on a single table (20%). Our real-world observations on LH workloads inspired this distribution. We execute the experiment on TPC-H sf1, sf10, and sf100 and report the average latency of all metadata operations of all runs in Fig. 9. Each phase lasted 20 minutes until a client switches to a new feature combination.
CAB-LV: Sequential Runs. Our hypothesis for sequential runs of read and write streams follows our previous observations. Read will benefit from configurations without added congestion with writes, while writes can use marker-based combinations to minimize the aborts. In Fig. 9(a), we show the latency of sequential metadata operations. We observe the best performance for LV [DL] and LV [I], that avoid the overhead of marker placement. In contrast, the worst performance is observed when we enable all LV features. LV [CT] ’s conflict detection has an overhead when enabled on LV [R] or LV [I]. On its own, LV [CT] has a high latency, but benefits from other features that use markers (compare LV [R, CT] with LV [CT, I]).
CAB-LV: Recommended Usage. We suggest using the original Delta Lake or LV [I] for reads, whereas for writes the decision depends on the desired isolation-level guarantees and the workload itself. Our recommendation is LV [DL] for single-query writes, LV [R] for single-table writes, and LV [R, CT] for multi-table writes.
CAB-LV: Concurrent Runs. We tested our recommendations by executing reads and writes concurrently in our CAB scenario. Fig. 9(b) shows a plot of the latencies of the metadata operations. In this setting, LV [I] exhibits the lowest latency due to its global version assignment. In contrast, with LV [DL] the readers experience a higher overhead because version assignments are done for all tables individually. Concurrent writes using LV [R] are unaffected by concurrent reads with LV [I], and can thus benefit of the marker placement. Effectively, when compared to LV [DL], LV [R] moves the main synchronization to the beginning of a transaction with a minor impact on the performance and stability using its sublog structure. When adding LV [CT] to both operations, we can see more outliers for the writes, and slower reads. The readers now need to set markers within the table, making them compete with the writers. However, by using LV [CT] we enable all clients to detect concurrent transactions and react to them (by adapting or requesting a new version) before committing. In summary, we confirm our recommendations, but we also note that the effects were only noticeable because we isolated the metadata operations. If we also account for data loading, then the observable effects are minimal.
5.5. Analytical workloads
Figure 10. TPC-DS: Data loading times of Spark and LakeVilla
[…]
[TPC-DS: Data loading times of Spark and LakeVilla]TPC-DS: Data loading times of Spark and LakeVilla. LakeVilla can retrieve the required data faster than Spark. Its features show minimal overhead.
Analytics and LakeVilla. We also analyze the impact of LV on OLAP using three TPC-DS queries from (Jain et al., 2023) on scale factors 1000 and 3000. These queries stress the system as they rely heavily on OTF metadata design but also require selected data (Jain et al., 2023). Our experiment includes LV [R], LV [CT], LV [I], Spark’s cold, and Spark’s best-of-three (hot) performance. We measure how long each client actively interacts with the object store using MinIO’s tracing features. Additionally, we investigate data caching effects and their influence on OTFs by comparing Spark’s hot and cold runs. We expect LV’s changes in the metadata layer operations will minimally impact performance as data loading via the network is the main bottleneck. While LV, as a storage layer, can not fully orchestrate query execution on its own or fully replicate Spark’s task concurrency, our results can be seen as a lower bound for full systems. Further, we expect noticeable differences between Spark’s hot and cold runs due to avoided object store loads.
Our results in Fig. 10 confirm our hypothesis. Queries like Q67 primarily depend on data loading, while Q68 and Q90 spend more time on execution-driven or intermediate result caching by Spark. Overall, since all LV features show similar performance for all queries, we conclude that LV’s features either do not or only exhibit a minor impact on OLAP performance.
5.6. Compatibility: LakeVilla + Existing Systems
(a) Estimation Spark Read
(b) Estimation Spark Write
Figure 11. Minimal additional requests needed for LakeVilla + Worst-case estimations for Delta Lake and Iceberg
[…]
[Minimal additional requests needed for LakeVilla + Worst-case estimations for Delta Lake and Iceberg.]Minimal additional requests needed for LakeVilla + Worst-case estimations for Delta Lake and Iceberg. Our estimations indicate that LakeVilla adds a minimal overhead in cost and possibly performance to Spark’s implementations.
Different Stages and Added Requests. However, one would argue that adding the extra complexity (features) not only induces overhead on the performance but also costs as we perform additional requests to the object store. To investigate, we measured the number of requests throughout our previous experiments and excluded requests not directed to OTF metadata layers. By comparing our baseline LV [DL] with the individual LV features, we determined the required requests for each feature. We expect LV to add a minor number of requests as its design integrates well into mandatory parts of the base OTFs.
Figure 11 shows the characteristics of Spark’s Delta Lake and Iceberg with the added requests for each feature for their single-query transactions. LV’s markers and sublog add the most additional write requests. For reads (Fig. 11(a)), this impact becomes more apparent as for writes (Fig. 11(b)), where Spark already issues such requests. Thus, this further confirms our previous results of LV [R] being more suitable for write transactions. LV [CT] adds a fixed number of one write request and five reads per accessed table. Hence, like its latency, its costs depend on the number of accessed tables within a transaction. Lastly, LV [I] adds the least amount of additional requests, with its added write requests being skippable for read-only transactions in its WriteSerializable mode, integrating it well into read transactions.
6. Related Work
Integrating and accessing data lakes. Previous research has focused on the operational efficiency of data lakes, like improving latency when accessing files by introducing strategies to encode relational data differently or improve file encodings (Kuschewski et al., 2023; Afroozeh and Boncz, 2023; Ghita et al., 2020). Further, Umbra focused on efficient parquet readers (Rey et al., 2023; Rey, 2024), transferring required sections of a parquet file and providing insights into the format structure. These approaches complement LV because Parquet files are commonly used for LH table data. Furthermore, cloud research has also explored elastic computing and additional structures to reduce cost and data lake traffic (Lee et al., 2023; Ritter, 2022; Vuppalapati et al., 2020; Winter et al., 2022; Durner et al., 2023; Perron et al., 2020; Müller et al., 2020; Kraft et al., 2023). These approaches follow the proposed architecture of external locks (Armbrust et al., 2020; Zaharia et al., 2021). In contrast, using a built-in approach, LV enables multiple features that previously required an auxiliary structure.
Lakehouses. Multiple proposed designs enabled missing features (Chen et al., 2022; Madden et al., 2022; Levandoski et al., 2024) or specializations (Begoli et al., 2021; Hambardzumyan, 2023) in custom OTFs. They redesign the metadata layer, introduce file optimizations and orderings, or add auxiliary structures and provide built-in features like security or performance for neural network training. Unlike these, LV is an extension applicable to existing OTFs. Project Nessie implements git-like branching for Apache Iceberg to isolate different transactions (cf. (i) & (iii)) (licence, 2024). However, Project Nessie is a client catalog, while LV focuses on built-in solutions. Similarly, Photon explores design decisions for a vectorized query engine processing Delta Lake tables like the combined memory management or row- versus column-oriented processing (Behm et al., 2022). Photon’s insights can become relevant for our LV prototype in the future as both use Delta Lake.
LakeVilla and External Solutions Multi-table transaction support is a heavily requested feature for LHs (Armbrust et al., 2020; Jain and Databricks, 2024; Foundation, 2024c, d), making it valuable to compare current advances to LV. The Delta Lake commit owner presumably is a hybrid solution, moving the commit decision to an external service and modifying Delta Lake to guarantee persistence (Jain and Databricks, 2024). In contrast, LV operates on the object store, making it a more independent solution. The possible advantages of the commit owners are better commit performance, tailored to Delta Lake, and cleaner metadata layers than LV’s markers. However, commit owners are presumably limited to a single vendor environment and not adaptive to other OTFs and transaction requirements. Additionally, the effects of an out-of-sync commit owner or the introduced ”backfilling” for compatibility are unknown and require further investigation (Jain and Databricks, 2024). We could see a combination of both used in practice: Commit Owners within a vendor’s environment for performance and LV for operations from the outside.
7. Conclusion
Current lakehouses provide limited transaction support without auxiliary structures. In this work, we introduce LakeVilla to improve transactional models and guarantees of state-of-the-art lakehouses. LakeVilla defines three modular features to achieve its goal and provides transactional guarantees up to serializability. Our results show the trade-offs and compatibility with systems used in practice. In contrast to other solutions in the Lakehouse area, we extend the base open table format on the object store non-invasively and allow transactions to freely choose their required guarantees, opening up interesting directions for future architectures and approaches.
Acknowledgements.
We are grateful to the colleagues at the Chair of Database Systems at TUM, the SAP HANA Cloud team, and all members of the SAP HANA Campus for their feedback. This work was partly funded by the Chair of Database Systems at TUM and SAP.